SPISAK OBJAVLJENIH NAUČNIH RADOVA

1. Pejović, N.: ”Preliminary analysis of the secular and main harmonic terms in the polar motion”, Publ. Dept. Astr. Univ. Beo- grad, 12, 1983, 41-48.

2. Pejović, N.: ”The stability the amplitudes and phases of the Chandler’s, annual and semi-annual term in the polar motion”, Publ. Dept. Astr. Univ. Beograd, 13, 1985, 31-39.

3. Pejović, N.: ”Short-period terms in the motion of the instantaneous pole”, Publ. Dept. Astr. Univ. Beograd, bf 14, 1986, 13-22.

4. Pejović, N.: ”Prilog izučavanju kretanja Zemljinih polova”, Univerzitet u Beogradu (magistarski rad), 1984.

5. Pejović, N.: ”Zemljina rotacija: posmatranja i atmosferska ekscitacija”, Saopštenje na IX Nacionalnoj konferenciji astronoma Jugoslavije, Sarajevo, 1988.

6. Vondrak, J.; Pejović, N.: ”Atmospheric excitation of the polar motion: comparison of polar motion spectrum with spectra of atospheric effective angular momentum functions”, Bull. Astron. Inst. Czechosl. 39, 1988, 172-185.

7. Pejović, N.; Vondrak, J.: ”Polar motion: observations and atmospheric excitation”, 6th International Symposium Geodesy and Physics of the Earth. Potsdam, 1988, Veroff. Zentralinst. Phys., Erde, 1989.

8. Pejović, N.: ”Globalna atmosferska cirkulacija i Zemljina rotacija”, Univerzitet u Beogradu (doktorska disertacija), 1989.

9. Pejović, N.; Vondrak,J.:”Atmospheric excitation of Earth’s rotation: comparison of the spectrum of length-of-day and axial component of the angular momentum function of the atmosphere”, Bull. Astron. Inst. Czechosl. 40, 1989, 382-393.

10. Pejović, N.: ”Earth’s model with variable Chandler’s frequency”, Bull. Astron. Inst. Czechosl. 41, 1990, 158-166.

11. Pejović, N.: ”Atmospheric excitation and variation in the length of day”, Publ. Dept. Astr. Univ. Beograd, 17, 1989.

12. Šegan, S.; Pejović, N.: ”The Motion of an Artificial Satellite in the Oblate Atmosphere”, Bull. Astron. Belgrade, 145, 1992, 101-102.

13. Pejović, N.; Šegan, S.: ”Frequency Modulation of the Čandler Wobble”, Bull. Astron. Belgrade, 145, 1992, 53-58.

14. Šegan, S.; Pejović, N.: ”Scale Heights of the Aeronomical Models and Perturbations of Satellite Orbits”, Bull. Astron. Belgrade, 148, 1993, 17-20.

15. Damljanović, G.; Pejovic, N:”Some results from the re-reduction of Belgrade zenithtelescope observations”, 1995 in Capitaine N., Kolaczek B., Debarbat S. (Ed), Proc. Earth Rotation, Reference System in Geodynamics and Solar System, Warsaw, Poland, 131- 133.

16. Damljanović, G; Pejović, N.: ”The Hipparcos mission and the Belgrade zenith-teles-cope observations”, Publ. Obs. Astron. Belgrade, 49, 1995, 127-130.

17. Damljanović, G.; Pejović, N.; -Durović, D.: ”The Chandler nutation for the period 1949-1985”. Bull. Astron. Belgrade, 156, 1997, 71-78.